# The asymptotic dimension of the grand arc graph is infinite

### v2

@@ Abstract.
Let $\Sigma$ be a compact, orientable surface, and let $\Gamma$ be a relation on $\pi_{0}(\partial\Sigma)$ such that the prescribed arc graph $\mathcal{A}(\Sigma,\Gamma)$ is Gromov-hyperbolic and non-trivial. We show that $\operatorname{asdim}\mathcal{A}(\Sigma,\Gamma)\geq-\chi(\Sigma)-1$, from which we prove that the asymptotic dimension of the grand arc graph is infinite. More generally, we prove that any connected, Gromov-hyperbolic multiarc and curve graph $\mathcal{M}$ preserved by $\operatorname{PMod}(\Sigma)$ with bounded geometric intersection over edges has $\operatorname{asdim}\mathcal{M}\geq g-\lceil\frac{1}{2}\chi(\Sigma)\rceil$, and that a broad class of multiarc and curve graphs on infinite-type surfaces has infinite asymptotic dimension.
@@ 1. Introduction
If $\mathcal{A}(\Sigma,\Gamma)$ is $\delta$-hyperbolic, then $-\chi(\Sigma)-1\leq\operatorname{asdim}\mathcal{A}(\Sigma,\Gamma).$
For $\Omega$ an infinite-type surface with finite grand splitting, let $\operatorname{\mathcal{G}}(\Omega)$ denote the grand arc graph on $\Omega$ [BNV22]. By applying Theorem 1.1, we obtain the following:
@@
If $\operatorname{\mathcal{G}}(\Omega)$ is non-empty and connected, then $\operatorname{asdim}\operatorname{\mathcal{G}}(\Omega)=\infty$.
More generally, let $\mathcal{M}$ be any connected multiarc and curve graph on a surface $\Omega$ that is preserved by $\operatorname{PMod}(\Omega)$, admits a (compact) witness, and in each witness has uniformly bounded geometric intersection over edges.
@@
In Section 2, we use the theory of alignment-preserving maps [DT17] to show that the Gromov boundary $\partial\mathcal{A}(\Sigma,\Gamma)$ contains $\partial\mathcal{A}(\Sigma)$. From results of Gabai [Gab14] and Schleimer [Po17] we obtain a compact subspace $Z\subset\partial\mathcal{A}(\Sigma,\Gamma)$ of dimension $-\chi(\Sigma)-2$. We then prove that $\operatorname{asdim}\mathcal{A}(\Sigma,\Gamma)\geq\dim Z+1$, extending a result for proper $\delta$-hyperbolic spaces, whence Theorem 1.1 follows.
In Section 3, we show that witness subsurfaces $W\subset\Omega$ for $\operatorname{\mathcal{G}}(\Omega)$ of arbitrarily large complexity admit prescribing relations $\Gamma$ such that $\mathcal{A}(W,\Gamma)$ quasi-isometrically embeds into $\operatorname{\mathcal{G}}(\Omega)$, where $\Omega$ is an infinite-type surface with finite grand splitting. In fact, $W$ may be chosen so that either $\mathcal{A}(W,\Gamma)$ has large coarse rank or it is $\delta$-hyperbolic: Theorem 1.2 thus follows from Theorem 1.1 and the monotonicity of asymptotic dimension.
Section 4 generalizes the techniques in Sections 2 and 3 to a broad class of simplicial graphs, called admissible combinatorial models, which include prescribed arc graphs, the grand arc graph, the $1$-skeleton of the marking complex, and many other multiarc and curve graphs. In addition to tools developed in Section 2, we utilize properties of the hierarchically hyperbolic structure of such graphs in the finite-type setting [Kop23a]. Theorem 1.3 follows from the analogous statements for admissible combinatorial models.
@@ 1.1. Background
An infinite-type surface $\Omega$ is classified by its genus and end space, which is obtained as the inverse limit of the complementary components of a compact exhaustion [Ric63]; its mapping class group $\operatorname{Mod}(\Omega)$ is a non-compactly generated Polish group. Given mild assumptions, Mann and Rafi [MR20] classify when $\operatorname{Mod}(\Omega)$ admits a generating set that is coarsely bounded (CB), or bounded in any left-invariant metric, and hence a well defined quasi-isometry type in the sense of [Ros14]. Mann–Rafi also define a preorder on the ends of $\Omega$ corresponding to topological complexity. We denote by $\mathscr{M}(\Omega)$ the non-empty subspace of maximal ends with respect to this preorder.
When $\operatorname{Mod}(\Omega)$ is locally CB (and in particular when it is CB-generated), Bar-Natan and Verberne define the grand splitting $\operatorname{\mathcal{S}}(\Omega)$, a canonical and $\operatorname{Mod}(\Omega)$-invariant partition of $\mathscr{M}(\Omega)$ into finitely many disjoint sets $E_{i}\in\operatorname{\mathcal{S}}(\Omega)$, each of which is either a singleton or Cantor set. A grand arc in $\Omega$ is a bi-infinite simple arc converging to ends in distinct sets in the grand splitting [BNV22].
@@
Let $\Omega$ be an infinite-type surface. The grand arc graph $\operatorname{\mathcal{G}}(\Omega)$ is the simplicial graph with vertices corresponding to isotopy classes of grand arcs and edges determined by disjointness.
The grand arc graph $\operatorname{\mathcal{G}}(\Omega)$ is a combinatorial model for $\Omega$ which generalizes the ray graph defined by Calegari [Cal09] on $S^{2}\setminus\text{Cantor set}$ and for surfaces with stable endspace extends the omnipresent arc graph defined by Fanoni–Ghaswala–McLeay [FGM21]. $\operatorname{Mod}(\Omega)$ acts naturally on $\operatorname{\mathcal{G}}(\Omega)$ by isometries. Bar-Natan–Verberne classify the $\delta$-hyperbolicity of $\operatorname{\mathcal{G}}(\Omega)$ and show that when $\operatorname{\mathcal{G}}(\Omega)$ is $\delta$-hyperbolic, the action of $\operatorname{Mod}(\Sigma)$ is quasi-continuous, extends continuously to $\partial\operatorname{\mathcal{G}}(\Omega)$, and has loxodromic elements.
@@ 1.1.1. Prescribed arc graphs, witnesses
Prescribed arc graphs were defined by the author in [Kop23b] as combinatorial models of finite-type surfaces that quasi-isometrically embed into $\operatorname{\mathcal{G}}(\Omega)$. Excepting trivial cases they are connected and infinite-diameter and their $\delta$-hyperbolicity is fully determined by the prescribing relation $\Gamma$:
@@
A compact, essential ($\pi_{1}$-injective, non-peripheral) non-pants subsurface is a witness for a given combinatorial model if it intersects every vertex. We call a witness $W\subset\Sigma$ for $\mathcal{A}(\Sigma,\Gamma)$ a $\Gamma$-witness.
@@ 1.1.2. Boundaries of non-proper $\delta$-hyperbolic spaces
In general, if $\mathcal{A}(\Sigma,\Gamma)$ is non-trivial then it is non-proper, and likewise for any admissible combinatorial model with sufficient complexity. For a geodesic $\delta$-hyperbolic space $X$, by $\partial X$ we always mean the sequential boundary of $X$; when $X$ is non-proper, $\partial X$ may be non-compact. In this setting, $\partial X$ does not coincide with the geodesic boundary, but is instead homeomorphic to the quasi-geodesic boundary [Has22]. We will make use of the following statement by Hasegawa, from a construction of Kapovich–Benakli [KB02, Rmk. 2.16]:
@@
Fixing $x_{0}\in X$, for any $z\in\partial X$ there exists a $(1+4\delta,12\delta)$-quasi-geodesic ray $\rho:[0,\infty)\to X$ based at $x_{0}$ with $[\rho(n)]=z$.
@@ 1.1.5. Alignment-preserving maps
We briefly recall the theory of alignment-preserving maps from [DT17]. Let $X$ be a geodesic metric space. Then a triple $(x,y,z)\in X^{3}$ is $K$-aligned if $d(x,y)+d(y,z)\leq d(x,z)+K$. A Lipschitz map between geodesic metric spaces $f:X\to Y$ is coarsely alignment preserving if there exists $K\geq 0$ for which $f$ maps any 00-aligned triple in $X$ to a $K$-aligned triple in $Y$.
@@ 2. Asymptotic dimension of $\mathcal{A}(\Sigma,\Gamma)$
By [Kop23b, §5], if $\Sigma\neq\Sigma_{0}^{4}$ and $\mathcal{A}(\Gamma,\Sigma)$ is $\delta$-hyperbolic, then (i) $\Sigma=\Sigma_{0}^{n+1}$and $\Gamma$ is an $n$-pointed star, (ii) $\Sigma=\Sigma_{1}^{2}$and $\Gamma$ is a non-loop edge, or (iii) $\Gamma$is not bipartite. In case (i), by [Kop23b, Lem. 5.4] $\mathcal{A}(\Sigma,\Gamma)$ is quasi-isometric to $\mathcal{A}(\Sigma,\ell_{0})$, where $\ell_{0}$ is a single loop and hence not bipartite. Thus for cases (i) and (iii), Corollary 2.2 implies $\partial\mathcal{A}(\Sigma)\subset\partial\mathcal{A}(\Sigma,\Gamma)$. In case (ii), every $\Gamma$-witness is in fact a witness for the usual arc graph: by [Kop23a] $\mathcal{A}(\Sigma,\Gamma)$ and $\mathcal{A}(\Sigma)$ have the same quasi-isometry type, hence $\partial\mathcal{A}(\Sigma,\Gamma)\cong\partial\mathcal{A}(\Sigma)$.
@@ 2.1. A lower bound
Applying the argument in [BL08, Prop. 6.5], we obtain that $\operatorname{asdim}\operatorname{Co}Z\geq\dim Z+1$. Proposition 2.5 then follows from Lemma 2.7. ///\!\!// /
3. Asymptotic dimension of $\operatorname{\mathcal{G}}(\Omega)$
@@
Since $\Sigma$ is a witness for $\operatorname{\mathcal{G}}(\Omega)$, it must separate distinct sets in $\operatorname{\mathcal{S}}(\Omega)$. In particular, each boundary component is adjacent to a complementary component containing ends in at most one set in $\operatorname{\mathcal{S}}(\Omega)$. Color each component $c\in\pi_{0}(\partial\Sigma)$ with the corresponding set $e(c)\in\operatorname{\mathcal{S}}(\Omega)$, if one exists; let $\Gamma$ be the complete $m$-partite relation on these colors (components without a corresponding class are left isolated).
@@
Fix a hyperbolic metric on $\Sigma$. For each colored boundary component $c$, choose a parameterization $c:[0,1)\to\Sigma$ and a simple ray $\rho_{c}$ disjoint from $\operatorname{int}(\Sigma)$ with origin $c(0)$ and converging to an end in $e(c)$. Let $a\in\mathcal{A}(\Sigma,\Gamma)$ be an arc that terminates on $c_{1},c_{2}\in\pi_{0}(\partial\Sigma)$. Let $\alpha$ be the geodesic representative for $a$ with endpoints $c_{i}(t_{i})$ and define $\delta_{i}=c_{i}|_{[0,t_{i}]}$ to be the subpath of $c_{i}$ between $c_{i}(0)$ and $c_{i}(t_{i})$. Let $\alpha^{\dagger}$ denote the extension of $\alpha$ from both endpoints by $\bar{\delta}_{i}*\rho_{c_{i}}$, for $i=1,2$ as appropriate. $\alpha^{\dagger}$ is a simple arc converging to ends in $e(c_{1}),e(c_{2})$ respectively, which are distinct in $\operatorname{\mathcal{S}}(\Omega)$ by our choice of $\Gamma$. $\alpha^{\dagger}$ is a grand arc. The map $a\mapsto[\alpha^{\dagger}]$ preserves disjointness hence extends to a simplicial ($1$-Lipschitz) map $\psi:\mathcal{A}(\Sigma,\Gamma)\to\operatorname{\mathcal{G}}(\Omega)$.
@@
We show that $\psi$ is a quasi-isometric embedding by constructing a coarse Lipschitz retraction $\pi:\operatorname{\mathcal{G}}(\Omega)\to\mathcal{A}(\Sigma,\Gamma)$. For a grand arc $w\in\operatorname{\mathcal{G}}(\Omega)$, fix a representative $\omega$ that is geodesic in $\Sigma$. Let $\omega^{\pm}$ denote the first and last intersections of $\omega$ with $\Sigma$ and let $\hat{\omega}$ denote the shortest path between $\omega^{-}$ and $\omega^{+}$ in $(\omega\cap\Sigma)\cup\partial\Sigma$. Since $\omega$ converges to maximal ends distinguished by $\operatorname{\mathcal{S}}(\Omega)$, $\omega^{\pm}$ lie on boundary components with distinct colors: isotoping $\hat{\omega}$ into the interior of $\Sigma$ rel $\omega^{\pm}$, $\hat{\omega}$ is $\Gamma$-allowed and we define $\pi:w\mapsto[\hat{\omega}]$. From the constructions of $\psi,\pi$, it is immediate that $\pi\psi$ is identity on $\mathcal{A}(\Sigma,\Gamma)$. We verify that $\pi$ is Lipschitz. Let $w,w^{\prime}\in\operatorname{\mathcal{G}}(\Omega)$ be disjoint grand arcs and let $\pi(w)=[\hat{\omega}]$ and $\pi(w^{\prime})=[\hat{\omega}^{\prime}]$ as above. Since $\hat{\omega}$ is constructed as a shortest path, it contains at most $|\pi_{0}(\partial\Sigma)|-1$ segments that are components of $\omega\cap\Sigma$. Each of these segments intersects $\hat{\omega}^{\prime}$ at most twice and in subsegments of $\hat{\omega}^{\prime}$ parallel to $\partial\Sigma$, and the same statement holds exchanging $\hat{\omega}$ and $\hat{\omega}^{\prime}$. Thus $i(\hat{\omega},\hat{\omega}^{\prime})\leq 4|\pi_{0}(\partial\Sigma)|-4$. Finally, since $d([\hat{\omega}],[\hat{\omega}^{\prime}])\leq i(\hat{\omega},\hat{\omega}^{\prime})+1$ by [Kop23b, Prop. 2.6], we obtain that $\pi$ is $(4|\pi_{0}(\partial\Sigma)|-3)$-Lipschitz. ∎
@@
Witnesses for $\operatorname{\mathcal{G}}(\Sigma)$ exist [BNV22, Lem. 2.7] and their enlargements are likewise witnesses, hence there exist fully separating witnesses $\Sigma\subset\Omega$ of arbitrarily large complexity. If $|\operatorname{\mathcal{S}}(\Omega)|>2$ and $\Gamma$ is chosen as in Lemma 3.2, then $\mathcal{A}(\Sigma,\Gamma)$ is $\delta$-hyperbolic by Theorem 1.5 and by Lemma 3.2 and Theorem 1.1 $\operatorname{asdim}\mathcal{G}(\Omega)>n$ for all $n$.
@@
Suppose instead that $|\operatorname{\mathcal{S}}(\Omega)|=2$. If $\Omega$ has infinite genus or infinitely many non-maximal ends, then there exists an infinite collection of pairwise-disjoint annular witnesses separating the sets $\{e,f\}=\operatorname{\mathcal{S}}(\Omega)$. Choosing finite subcollections defines quasi-flats of arbitrarily large dimension [Sch, Exercise 3.13], hence again $\operatorname{asdim}\operatorname{\mathcal{G}}(\Omega)=\infty$. Alternatively, $\operatorname{\mathcal{G}}(\Omega)$ contains an asymphoric hierarhically hyperbolic space of arbitrarily high rank, hence has infinite asymptotic dimension [Kop23a, Prop. 1.11].
@@
Finally, suppose that $|\operatorname{\mathcal{S}}(\Omega)|=2$ and $\Omega$ has finite genus and finitely many non-maximal ends. $\Omega$ must have at least one infinite set $e\in\operatorname{\mathcal{S}}(\Omega)$; let $f\in\operatorname{\mathcal{S}}(\Omega)$ be the other set. For any $n$, choose a $(n+1)$-holed sphere $\Sigma\subset\Omega$ with $n$ boundary components partitioning $e$ and the remaining component separating $e$ from $f$ and any genus or non-maximal ends. Then $\Sigma$ is a fully separating witness for $\operatorname{\mathcal{G}}(\Omega)$ and $\Gamma$, defined as in Lemma 3.2, is a $n$-pointed star. $\mathcal{A}(\Sigma,\Gamma)$ is $\delta$-hyperbolic by Theorem 1.5: we conclude by Lemma 3.2 and Theorem 1.1. ///\!\!// /
4. Asymptotic dimension of arbitrary combinatorial models
We generalize the preceding arguments to a broad class of combinatorial models for finite and infinite-type surfaces.
4.1. Admissible combinatorial models
Let $\Omega$ be a orientable surface of finite or infinite type. We first provide an extension of arc and curve systems and markings on $\Omega$ that subsumes both.

### v3

@@ Abstract.
Let $\Sigma$ be a compact, orientable surface of genus $g$, and let $\Gamma$ be a relation on $\pi_{0}(\partial\Sigma)$ such that the prescribed arc graph $\mathcal{A}(\Sigma,\Gamma)$ is Gromov-hyperbolic and non-trivial. We show that $\operatorname{asdim}\mathcal{A}(\Sigma,\Gamma)\geq-\chi(\Sigma)-1$, from which we prove that the asymptotic dimension of the grand arc graph is infinite. More generally, an arc and curve model on $\Sigma$ is a graph of simple arc and curves on $\Sigma$, on which $\operatorname{PMap}(\Sigma)$ acts by permuting vertices. We prove that any connected, Gromov-hyperbolic cocompact arc and curve model $\mathcal{M}$ has $\operatorname{asdim}\mathcal{M}\geq g-\lceil\frac{1}{2}\chi(\Sigma)\rceil$, and that a broad class of arc and curve models on infinite-type surfaces has infinite asymptotic dimension.
@@ 1. Introduction
If $\mathcal{A}(\Sigma,\Gamma)$ is $\delta$-hyperbolic, then $\chi(\Sigma)-1\leq\operatorname{asdim}\mathcal{A}(\Sigma,\Gamma).$
For $\Omega$ an infinite-type surface with finite grand splitting, let $\mathcal{G}(\Omega)$ denote the grand arc graph on $\Omega$ [BNV22]. By applying Theorem 1.1, we obtain the following:
@@
If $\mathcal{G}(\Omega)$ is non-empty and connected, then $\operatorname{asdim}\mathcal{G}(\Omega)=\infty$.
More generally, an arc and curve model on a surface $\Omega$ is a connected graph whose vertices are collections of (possibly intersecting) simple arcs and curves, with an action of $\operatorname{PMap}_{c}(\Omega)$ induced by the permutation of its vertices. An arc and curve model $\mathcal{M}$ is witness-cocompact if it has a (compact) witness and in each witness has uniformly bounded geometric (self-)intersection over the vertices and edges in $\mathcal{M}$.
@@
In Section 2, we use the theory of alignment-preserving maps [DT17] to show that the Gromov boundary $\partial\mathcal{A}(\Sigma,\Gamma)$ contains $\partial\mathcal{A}(\Sigma)$. From results of Gabai [Gab14] and Schleimer [Po17] we obtain a compact subspace $Z\subset\partial\mathcal{A}(\Sigma,\Gamma)$ of dimension $\chi(\Sigma)-2$. We then prove that $\operatorname{asdim}\mathcal{A}(\Sigma,\Gamma)\geq\dim Z+1$, extending a result for proper $\delta$-hyperbolic spaces, whence Theorem 1.1 follows.
In Section 3, we show that witness subsurfaces $W\subset\Omega$ for $\mathcal{G}(\Omega)$ of arbitrarily large complexity admit prescribing relations $\Gamma$ such that $\mathcal{A}(W,\Gamma)$ quasi-isometrically embeds into $\mathcal{G}(\Omega)$, where $\Omega$ is an infinite-type surface with finite grand splitting. In fact, $W$ may be chosen so that either $\mathcal{A}(W,\Gamma)$ has large coarse rank or it is $\delta$-hyperbolic: Theorem 1.2 thus follows from Theorem 1.1 and the monotonicity of asymptotic dimension.
Section 4 generalizes the techniques in Sections 2 and 3 to witness-cocompact arc and curve models, which include prescribed arc graphs, the grand arc graph, the marking complex, and many other multiarc and curve graphs. In addition to tools developed in Section 2, we utilize properties of the hierarchically hyperbolic structure of such graphs in the finite-type setting [Kop23a]. Theorem 1.3 follows analogously in the witness-cocompact case.
@@ 1.1. Background
An infinite-type surface $\Omega$ is classified by its genus and end space, which is obtained as the inverse limit of the complementary components of a compact exhaustion [Ric63]; its mapping class group $\operatorname{Map}(\Omega)$ is a non-compactly generated Polish group. Given mild assumptions, Mann and Rafi [MR20] classify when $\operatorname{Map}(\Omega)$ admits a generating set that is coarsely bounded (CB), or bounded in any left-invariant metric, and hence a well defined quasi-isometry type in the sense of [Ros14]. Mann–Rafi also define a preorder on the ends of $\Omega$ corresponding to topological complexity. We denote by $\mathscr{M}(\Omega)$ the non-empty subspace of maximal ends with respect to this preorder.
When $\operatorname{Map}(\Omega)$ is locally CB (and in particular when it is CB-generated), Bar-Natan and Verberne define the grand splitting $\operatorname{\mathcal{S}}(\Omega)$, a canonical and $\operatorname{Map}(\Omega)$-invariant partition of $\mathscr{M}(\Omega)$ into finitely many disjoint sets $E_{i}\in\operatorname{\mathcal{S}}(\Omega)$, each of which is either a singleton or Cantor set. A grand arc in $\Omega$ is a bi-infinite simple arc converging to ends in distinct sets in the grand splitting [BNV22].
@@
Let $\Omega$ be an infinite-type surface. The grand arc graph $\mathcal{G}(\Omega)$ is the simplicial graph with vertices corresponding to isotopy classes of grand arcs and edges determined by disjointness.
The grand arc graph $\mathcal{G}(\Omega)$ is an arc and curve model for $\Omega$ which generalizes the ray graph defined by Calegari [Cal09] on $S^{2}\setminus\text{Cantor set}$ and for surfaces with stable endspace extends the omnipresent arc graph defined by Fanoni–Ghaswala–McLeay [FGM21]. $\operatorname{Map}(\Omega)$ acts naturally on $\mathcal{G}(\Omega)$ by isometries. Bar-Natan–Verberne classify the $\delta$-hyperbolicity of $\mathcal{G}(\Omega)$ and show that when $\mathcal{G}(\Omega)$ is $\delta$-hyperbolic, the action of $\operatorname{Map}(\Sigma)$ is quasi-continuous, extends continuously to $\partial\mathcal{G}(\Omega)$, and has loxodromic elements.
@@ 1.1.1. Prescribed arc graphs, witnesses
Prescribed arc graphs were defined by the author in [Kop23b] as arc and curve models of finite-type surfaces that quasi-isometrically embed into $\mathcal{G}(\Omega)$. Excepting trivial cases they are connected and infinite-diameter and their $\delta$-hyperbolicity is fully determined by the prescribing relation $\Gamma$:
@@
A compact, essential ($\pi_{1}$-injective, non-peripheral) subsurface without pants components is a witness for a given arc and curve model if each component intersects every vertex. We call a witness $W\subset\Sigma$ for $\mathcal{A}(\Sigma,\Gamma)$ a $\Gamma$-witness.
@@ 1.1.2. Boundaries of non-proper $\delta$-hyperbolic spaces
In general, if $\mathcal{A}(\Sigma,\Gamma)$ is non-trivial then it is non-proper, and likewise for any admissible arc and curve model with sufficient complexity. For a geodesic $\delta$-hyperbolic space $X$, by $\partial X$ we always mean the sequential boundary of $X$; when $X$ is non-proper, $\partial X$ may be non-compact. In this setting, $\partial X$ does not coincide with the geodesic boundary, but is instead homeomorphic to the quasi-geodesic boundary [Has22]. We will make use of the following statement by Hasegawa, from a construction of Kapovich–Benakli [KB02, Rmk. 2.16]:
@@
Fixing $x_{0}\in X$, for any $z\in\partial X$ there exists a $(1+4\delta,12\delta)$-quasi-geodesic ray $\rho:[0,\infty)\to X$ based at $x_{0}$ with $\rho(n)]=z$.
@@ 1.1.4. Markings
Definition 1.11 ([Kop23a, Def. 2.2]).
A marking $\mu$ on a surface $\Omega$ is locally clean if the maximal submarking $\mu^{\prime}$ with only non-empty transversals is complete and clean in each component of $\Omega\setminus(\mu\setminus\mu^{\prime})$ that it intersects.
@@ 1.1.5. Alignment-preserving maps
We briefly recall the theory of alignment-preserving maps from [DT17]. Let $X$ be a geodesic metric space. Then a triple $(x,y,z)\in X^{3}$ is $K$-aligned if $d(x,y)+d(y,z)\leq d(x,z)+K$. A Lipschitz map between geodesic metric spaces $f:X\to Y$ is coarsely alignment preserving if there exists $K\geq 0$ for which $f$ maps any 0-aligned triple in $X$ to a $K$-aligned triple in $Y$.
@@ 2. Asymptotic dimension of $\mathcal{A}(\Sigma,\Gamma)$
[…]
@@ 2.1. A lower bound
Applying the argument in [BL08, Prop. 6.5], we obtain that $\operatorname{asdim}\operatorname{Co}Z\geq\dim Z+1$. Proposition 2.5 then follows from Lemma 2.7. ///\!\!/
3. Asymptotic dimension of $\mathcal{G}(\Omega)$
@@
Since $\Sigma$ is a witness for $\mathcal{G}(\Omega)$, it must separate distinct sets in $\operatorname{\mathcal{S}}(\Omega)$. In particular, each boundary component is adjacent to a complementary component containing ends in at most one set in $\operatorname{\mathcal{S}}(\Omega)$. Color each component $c\in\pi_{0}(\partial\Sigma)$ with the corresponding set $e(c)\in\operatorname{\mathcal{S}}(\Omega)$, if one exists; let $\Gamma$ be the complete $m$-partite relation on these colors (components without a corresponding class are left isolated).
@@
Fix a hyperbolic metric on $\Sigma$. For each colored boundary component $c$, choose a parameterization $c:[0,1)\to\Sigma$ and a simple ray $\rho_{c}$ disjoint from $\operatorname{int}(\Sigma)$ with origin $c(0)$ and converging to an end in $e(c)$. Let $a\in\mathcal{A}(\Sigma,\Gamma)$ be an arc that terminates on $c_{1},c_{2}\in\pi_{0}(\partial\Sigma)$. Let $\alpha$ be the geodesic representative for $a$ with endpoints $c_{i}(t_{i})$ and define $\delta_{i}=c_{i}|_{[0,t_{i}]}$ to be the subpath of $c_{i}$ between $c_{i}(0)$ and $c_{i}(t_{i})$. Let $\alpha^{\dagger}$ denote the extension of $\alpha$ from both endpoints by $\bar{\delta}_{i}*\rho_{c_{i}}$, for $i=1,2$ as appropriate. $\alpha^{\dagger}$ is a simple arc converging to ends in $e(c_{1}),e(c_{2})$ respectively, which are distinct in $\operatorname{\mathcal{S}}(\Omega)$ by our choice of $\Gamma$. $\alpha^{\dagger}$ is a grand arc. The map $a\mapsto[\alpha^{\dagger}]$ preserves disjointness hence extends to a simplicial ($1$-Lipschitz) map $\psi:\mathcal{A}(\Sigma,\Gamma)\to\mathcal{G}(\Omega)$.
@@
We show that $\psi$ is a quasi-isometric embedding by constructing a coarse Lipschitz retraction $\pi:\mathcal{G}(\Omega)\to\mathcal{A}(\Sigma,\Gamma)$. For a grand arc $w\in\mathcal{G}(\Omega)$, fix a representative $\omega$ that is geodesic in $\Sigma$. Let $\omega^{\pm}$ denote the first and last intersections of $\omega$ with $\Sigma$ and let $\hat{\omega}$ denote the shortest path between $\omega^{-}$ and $\omega^{+}$ in $(\omega\cap\Sigma)\cup\partial\Sigma$. Since $\omega$ converges to maximal ends distinguished by $\operatorname{\mathcal{S}}(\Omega)$, $\omega^{\pm}$ lie on boundary components with distinct colors: isotoping $\hat{\omega}$ into the interior of $\Sigma$ rel $\omega^{\pm}$, $\hat{\omega}$ is $\Gamma$-allowed and we define $\pi:w\mapsto[\hat{\omega}$]. From the constructions of $\psi,\pi$, it is immediate that $\pi\psi$ is identity on $\mathcal{A}(\Sigma,\Gamma)$. We verify that $\pi$ is Lipschitz. Let $w,w^{\prime}\in\mathcal{G}(\Omega)$ be disjoint grand arcs and let $\pi(w)=[\hat{\omega}$] and $\pi(w^{\prime})=[\hat{\omega}^{\prime}]$ as above. Since $\hat{\omega}$ is constructed as a shortest path, it contains at most $|\pi_{0}(\partial\Sigma)|-1$ segments that are components of $\omega\cap\Sigma$. Each of these segments intersects $\hat{\omega}^{\prime}$ at most twice and in subsegments of $\hat{\omega}^{\prime}$ parallel to $\partial\Sigma$, and the same statement holds exchanging $\hat{\omega}$ and $\hat{\omega}^{\prime}$. Thus $i(\hat{\omega},\hat{\omega}^{\prime})\leq 4|\pi_{0}(\partial\Sigma)|-4$. Finally, since $d([\hat{\omega}],[\hat{\omega}^{\prime}])\leq i(\hat{\omega},\hat{\omega}^{\prime})+1$ by [Kop23b, Prop. 2.6], we obtain that $\pi$ is $(4|\pi_{0}(\partial\Sigma)|-3)$-Lipschitz. ∎
@@
Witnesses for $\mathcal{G}(\Sigma)$ exist [BNV22, Lem. 2.7] and their enlargements are likewise witnesses, hence there exist fully separating witnesses $\Sigma\subset\Omega$ of arbitrarily large complexity. If $|\operatorname{\mathcal{S}}(\Omega)|>2$ and $\Gamma$ is chosen as in Lemma 3.2, then $\mathcal{A}(\Sigma,\Gamma)$ is $\delta$-hyperbolic by Theorem 1.5 and by Lemma 3.2 and Theorem 1.1 $\operatorname{asdim}\mathcal{G}(\Omega)>n$ for all $n$.
@@
Suppose instead that $|\operatorname{\mathcal{S}}(\Omega)|=2$. If $\Omega$ has infinite genus or infinitely many non-maximal ends, then there exists an infinite collection of pairwise-disjoint annular witnesses separating the sets $\{e,f\}=\operatorname{\mathcal{S}}(\Omega)$. Choosing finite subcollections defines quasi-flats of arbitrarily large dimension [Sch, Exercise 3.13], hence again $\operatorname{asdim}\mathcal{G}(\Omega)=\infty$. Alternatively, $\mathcal{G}(\Omega)$ contains an asymphoric hierarhically hyperbolic space of arbitrarily high rank, hence has infinite asymptotic dimension [Kop23a, Prop. 1.11].
@@
Finally, suppose that $|\operatorname{\mathcal{S}}(\Omega)|=2$ and $\Omega$ has finite genus and finitely many non-maximal ends. $\Omega$ must have at least one infinite set $e\in\operatorname{\mathcal{S}}(\Omega)$; let $f\in\operatorname{\mathcal{S}}(\Omega)$ be the other set. For any $n$, choose a $(n+1)$-holed sphere $\Sigma\subset\Omega$ with $n$ boundary components partitioning $e$ and the remaining component separating $e$ from $f$ and any genus or non-maximal ends. Then $\Sigma$ is a fully separating witness for $\mathcal{G}(\Omega)$ and $\Gamma$, defined as in Lemma 3.2, is a $n$-pointed star. $\mathcal{A}(\Sigma,\Gamma)$ is $\delta$-hyperbolic by Theorem 1.5: we conclude by Lemma 3.2 and Theorem 1.1. ///\!\!/
4. Asymptotic dimension of arc and curve models
We generalize the preceding arguments to a broad class of arc and curve models for finite and infinite-type surfaces. We first consider cocompact arc and curve models, which we demonstrate to be equivariantly quasi-isometric to cocompact marking graphs in the finite-type case; it follows that such arc and curve models are hierarchically hyperbolic by [Kop23a]. We then compute asymptotic dimension lower bounds for such models, which we use to construct lower bounds in the infinite-type setting.
4.1. Cocompact arc and curve models
Let $\Omega$ be a connected and non-pants hyperbolic surface of finite or infinite type. We first provide an extension of arc and curve systems and markings on $\Omega$ that subsumes both. Let $\mathcal{K}(\Omega)\mathrel{\mathop{:}}=K(V(\mathcal{AC}(\Omega)))$ denote the set of finite collections of (not necessarily disjoint) simple arcs and curves on $\Omega$ and let $\operatorname{PMap}_{c}(\Omega)\leq\operatorname{Map}(\Omega)$ denote the subgroup of compactly supported pure mapping classes.